### v1

# State of practice: evaluating GPU performance of state vector and tensor network methods

Abstract
The frontier of quantum computing (QC) simulation on classical hardware is quickly reaching the hard scalability limits for computational feasibility. Nonetheless, there is still a need to simulate large quantum systems classically, as the Noisy Intermediate Scale Quantum (NISQ) devices are yet to be considered fault tolerant and performant enough in terms of operations per second. Each of the two main exact simulation techniques, state vector and tensor network simulators, boasts specific limitations. The exponential memory requirement of state vector simulation, when compared to the qubit register sizes of currently available quantum computers, quickly saturates the capacity of the top HPC machines currently available. Tensor network contraction approaches, which encode quantum circuits into tensor networks and then contract them over an output bit string to obtain its probability amplitude, still fall short of the inherent complexity of finding an optimal contraction path, which maps to a max-cut problem on a dense mesh, a notably NP-hard problem.
This article aims at investigating the limits of current state-of-the-art simulation techniques on a test bench made of eight widely used quantum subroutines, each in 31 different configurations, with special emphasis on performance. We then correlate the performance measures of the simulators with the metrics that characterise the benchmark circuits, identifying the main reasons behind the observed performance trend. From our observations, given the structure of a quantum circuit and the number of qubits, we highlight how to select the best simulation strategy, obtaining a speedup of up to an order of magnitude.
Index Terms:
Quantum, Tensor Network, HPC
I Introduction
Do we really need to build quantum computers?
The answer is blatantly affirmative, however knowing where the hard boundary for the classical simulation of quantum systems lies is most definitely not a straightforward task.
The quantum computing paradigm, since its theoretical inception by one of the forefathers of modern physics R. Feynman in 1982[14], served to extend the classical definition of computation in order to better describe the quantum properties of nature, under the hypothesis that an experiment, which is a purposefully built physical system, can be said to be performing computations under specific conditions[20]. Years later, many quantum algorithms would be proposed, able to exploit the characteristics of this new paradigm to achieve unprecedented speedups in unstructured search problems [17] or to efficiently solve hard mathematical problems, such as the prime factorization of large numbers [28], and others in the fields of optimisation, machine learning and chemistry.
In the last decade, the progress in the various technologies used for building quantum devices has reached commercial applications, such as the cloud services proposed by IBM, D-Wave and Google, among others. These systems, however, are yet to be considered mature, as most still fail in preserving coherent quantum states, suffer cross-talk among their constituent elements and employ imperfect logical operators. Furthermore, supercondution-based devices, which make up a wide margin of the operational quantum devices as of the writing of this article, have been recently proven to be exceedingly susceptible to external radiation events [35, 37, 24], once again hindering their applicability.
Meanwhile, the top performing supercomputers have surpassed the exascale number of floating operations per second [32]. Naturally, such an achievement pushes up the threshold for the complexity of quantum systems that can be classically simulated over a reasonable time span [19]. This opens up the possibility for some quantum algorithms to be efficiently converted into quantum-inspired classical algorithms. Moreover, the testing and validation of newly proposed quantum algorithms, and the accuracy measurement of real quantum devices’ outputs must be done through simulation, to double check results and ensure correct operation. With the increasingly higher optimisation of specialised software libraries over commodity hardware accelerators, such as graphic processing units (GPUs), simulation of small quantum algorithms is becoming more and more easy to accomplish [29, 30, 1, 5]. Such practices are, of course, still limited in memory and time by the exponential requirements of quantum simulation, but better exploitation of current classical computation resources may lead to efficient simulation of small quantum subroutines without needing to compensate for hardware-level noise and external radiation events, possibly reducing or annulling the queries made to cloud-based quantum computers for low qubit sized problems.
The purpose of this article is to understand where the limit for efficient quantum simulation on classical hardware lies, emphasising the computational aspects, including the time and memory footprints of quantum algorithms. There are various classical simulation techniques for quantum algorithms other than state vector simulation and tensor network contractions, such as stabilisers theory or p-block simulation. Notably, however, stabiliser theory is limited to the application of a non-universal set of Clifford gates, whilst p-block simulation leverages approximations in the representation that lead to very limited entanglement. Being the only two approaches able to find an exact solution, the analysis we propose focuses on state vector simulation and tensor network contraction techniques.
We prove that, by profiling quantum circuits with the approach presented in this paper, the simulation time can reach a speedup of up to one order of magnitude, especially for large quantum circuits. The proposed performance-circuit characteristics correlation is achieved by characterising a purposefully selected suite of well known quantum circuit subroutines according to objective metrics, and checking how those scale with respect to the size of the quantum circuit. All the circuits we consider are parameterisable over the number of qubits in the system, and some of them feature additional customisation parameters, such as layer repetition. Moreover, they have been selected as to have practical applicability in terms of exact simulation. These same circuits will be then simulated on top level hardware, using both state-vector and tensor network contraction approaches through NVIDIA’s cuQuantum library [29], highlighting which one boasts the better performance for each workload. Our work proposes a practical methodology to pick the most efficient simulation strategy according to a given set of static characteristics of the circuit.
The rest of the paper is organized as follows: Section II provides a short summary of quantum computing as a whole, and it is followed by Section III, which introduces how classical algorithms for quantum simulation work. Section IV gives a definition of the metrics and of the quantum circuits considered for this study. Section V characterises the quantum circuits according to the aforementioned metrics, then presents performance results with respect to execution times and peak memory occupancy. Lastly, Section VI concludes the paper by expanding on the hereby presented work by opening new paths for investigation in future works.
II Background
Nature, on an atomic and subatomic scale, is inherently quantum. When tasked with modeling and simulating the properties of such atomic-scale phenomena, it is reasonable to do so with objects that are able to express the same quantum properties that are to be investigated. This implies that the binary computation paradigm has to evolve towards a representation which encompasses such additional characteristics.
II-A Quantum Computing
Quantum computing is an expansion of binary computing able to tackle any problem that the latter approach can tackle, whilst at the same time providing more efficient solutions to problems that are deemed intractable in the classical domain. This is achieved by exploiting ad hoc resources, namely superposition and entanglement.
Quantum computers make use of qubits, the quantum counterpart of classical bits, to encode information. Each qubit is described by two complex probability amplitudes, as follows
[…]
Each amplitude, when squared, represents the probability for the qubit to collapse on its corresponding classical state when observed. It must hold for the sum of the two squared probabilities to equal unity, i.e. $\alpha^{2}+\beta^{2}=1$. Whenever the two amplitudes assume non-integer values, the qubit is said to be in superposition between the two basis states. The property of entanglement refers to the fact that two or more qubits can share a non-classical correlation, such that when one of the entangled qubits is observed, it lets us infer information regarding the state of the others without measuring them. This second property sprouted issues with locality of quantum mechanics in the well-known EPR paradox [11], later solved by J.S. Bell [4], which defined the minimum set of these namesake entangled quantum states. These two qubit states cannot be described as the product of two independent qubits, such as
[…]
Entanglement is believed to be the fundamental resource responsible for quantum speedup, although superposition plays an important role as well, since circuits with low entanglement have been proven to be easy to simulate [36].
II-B Quantum circuits
Algorithms in the QC field are expressed via quantum circuits, a graphical notation derived from Penrose’s notation[9]. They are read from left to right, following the flow of information. Operations on one or multiple qubits are applied via quantum gates, which are represented by $2^{N}\times 2^{N}$ unitary matrices, with $N$ being the number of qubits acted upon by the gate. The most basic single qubit operators include the Pauli X, Y, Z gates and the basis-swap Hadamard gate, represented by the following matrices
$\begin{split}X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},&\quad Y=\begin{bmatrix}0&-i\\ i&0\end{bmatrix}\\ Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},&\quad H=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}\end{split}$ (3)
[…]
$CNOT=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{bmatrix}$ (4)
Similarly to the binary paradigm, it is possible to define universal quantum gate sets that can be used to encode any operator as the composition of these fundamental ones. This is at the basis of the construction of universal quantum computers.
[…]
Figure 1: Comparison between the theoretical output distribution of the Bell circuit (left) and the experimental output distribution obtained from a real quantum device, an IBM Falcon r4T processor, over 1024 measurement shots (right).
[…]
The execution of a quantum circuit on a real quantum device yields a single classical output bit string. By repeatedly executing and measuring the same circuit, it is possible to sample an output distribution, which converges to the element-wise square of the state vector made of the probability amplitudes of each possible quantum state in the system. Experimental results and expected theoretical results oftentimes differ, due to intrinsic errors in the quantum device, cross talk among qubits and the number of samples made. Such an example has been provided in Figure 1. Classical simulation of quantum systems is mainly used to compute the expected theoretical output distribution of a quantum circuit, so as to later compare it to the output of a real quantum device.
III Quantum circuit simulation
Quantum simulation can refer to two concepts: either the usage of real quantum devices to simulate other quantum systems, or the usage of classical machines to compute the theoretical output of a quantum algorithm. For the sake of clarity, this paper will always refer to the latter when talking about quantum circuit simulation. The objective of simulation is not to thwart the development of real quantum devices, but rather to validate the outputs of such machines against their theoretical expected outputs. Moreover, given the still relatively scarce availability of real quantum devices and the limitations of current device technology, such as low coherence times, quantum simulators provide a means for validating new and possibly deeper quantum algorithms. There are various approaches to simulate a quantum circuit, the two main ones that provide exact results are: state vector simulation [12] and tensor network contraction [16], which are detailed in the following subsections.
We aim at understanding how the inherent structure of a quantum circuit can affect the execution time, so as to preemptively identify which simulation strategy works best for which kind of circuit, by making use of ad hoc metrics. These metrics provide a description of the overall structure of the quantum circuit, highlighting critical areas for the improvement of modern simulators. It will be possible to infer that any other quantum circuit, reflecting the characteristics provided in this paper, will scale similarly in terms of simulation. The main current limitation of state vector simulators is the inherent exponential memory blowout linked to the system size, that has been tentatively compensated through state vector compression [38], however the distributed application of vector-matrix multiplications still scales exponentially on the system size. Tensor networks have already shown promising results, with useful applications in the field of verification of real quantum computer’s outputs, however due to the limited exploitation of the internal structures of the circuit-derived graph representation, the contraction path used is not granted to be optimal.
III-A State vector simulation
Quantum states are represented by a wave function, which can be encoded into a state vector. Given any quantum system of $N$ qubits, its corresponding state vector will contain $2^{N}$ complex probability amplitudes, one for each possible output bit string. Quantum gates are applied by splitting the state vector into smaller vectors of size equal to that of the gate to be applied, then each sub-vector is multiplied with the gate matrix and the resulting sub-vectors are reassembled in the evolved state vector. The splitting operation is performed according to the qubits over which the operator is applied. This can be intuitively understood by considering the ordered set of output bit strings: the probability amplitudes corresponding to a given sequence of qubits, which depends on the qubit indices the operator acts onto, are grouped together. An example of this process for both single and two qubit gates is depicted in Figure 3.
Figure 3: Splitting of the state vector and application of the vector-matrix multiplication in a 3 qubit system. On the left, a single qubit gate is applied to qubit 1, so the amplitude pairs are grouped by following a 0-1 repeated scheme for the amplitude’s index. On the right, a double qubit gate is applied on qubits 0 and 1, so the amplitude pairs are grouped following a 00-01-10-11 scheme. All amplitude indices are written in little endian and ordered top to bottom.
Figure 4: The process of converting a quantum circuit into a tensor network to extract a probability amplitude. On the left, an example of a quantum circuit. In the center, the circuit gets converted into a tensor network representation, where single and double qubit gates become order-2 and order-3 tensors, respectively. On the right, after the tensor network contraction, we get the probability amplitude of a specific bit string: by repeating this process over all output bit strings, one may reconstruct the whole state vector.
The state vector simulation’s complexity scales linearly in time with respect to the number of gates [12]. However, the memory footprint of the state vector and the number of vector-matrix multiplications performed increase exponentially with the number of qubits present in the system to be simulated.
[…]
To put that into perspective, it is possible to roughly estimate the number of atoms in the observable universe to be $10^{82}\approx 2^{270}$ [27]: this means that, if we were to to store a single amplitude value inside each of them to represent a state vector, we could only represent state vectors of systems with up to 270 qubits. Well known quantum algorithms need significantly more logical qubits [18], and this is without considering the cost in terms of classical computation time, which may add up to reach unfathomable time scales[3]. Overall, the state vector approach is generally convenient when simulating small quantum systems, as it produces a full description of the output wavefunction.
III-B Tensor network simulation
Quantum gates and quantum basis states are represented by tensors. The graphical representation of a quantum circuit can be read as a directed acyclic graph (DAG), where the vertices are represented by quantum gates or basis states and the edges are represented by the qubit wires. The input tensors are the basis states, encoded as follows
[…]
All other gates use their standard matrix representation. The contraction of an edge corresponds to the multi-dimensional generalisation of the dot product between two tensors over a shared index. The measurement operators at the end of the quantum circuit are substituted by open indices. Whenever a full network contraction operation is performed, the output open indices are closed with the conjugate tensors of the basis states that represent a specific bit string. Doing so, followed by the contraction of the newly closed indices, produces the probability amplitude of the chosen bit string. Figure 4 provides graphical insight over this process.
It is possible to use tensor networks to reconstruct the whole state vector, by closing and contracting the output indices over different bit strings, however doing so incurs in the same limits of the state vector simulator for storing the final vector.
[…]
The memory occupancy of the tensor network grows linearly with respect to the number of quantum gates and qubits in the system. This approach moves the complexity of simulation to that of finding an optimal contraction path for the tensor network, which is known to be an NP-hard problem [15]. Efficient heuristics specialised for quantum circuit-derived tensor networks have been proposed [16], however there is no catch all heuristic for this kind of problem in general. If the contraction path is not optimal, it may lead to increased computation time, possibly making it less efficient than state vector simulation altogether. To the interested reader, we suggest some resources for tensor network theory by J. Biamonte [8, 7].
IV Benchmarks and metrics
To assess the performance of current state of the art simulators and to select the most efficient one, it is necessary to use a set of metrics and quantum circuits which are relevant and well established in the quantum computing field. We rely on two quantum circuit benchmarking suites, which are widely recognised in the literature: SupermarQ [31] and QASMBench [23]. Both of these suites provide their own sets of quantum circuits, that, however, have been specifically selected for testing the hardware performance of real quantum devices. For this reason, some of these circuits boast little to no practical use in the context of noiseless exact simulation, such as the error correcting code circuits in SupermarQ, or the Greenberger–Horne–Zeilinger.
[…]
Furthermore, both suites introduce a list of metrics that characterise the topological nature of static quantum circuits. These metrics provide a measure of topological properties of the graph derived from the quantum circuit representation, letting us correlate such properties with the runtime performance statistics.
IV-A SupermarQ
In the SupermarQ [31] suite, six metrics are introduced, however we will only consider the ones that have topological significance, referring to all elements which may alter the circuit-derived graph, that is the relative presence and the disposition of two-qubit or higher size quantum gates. All metrics range in $[0,1]$, where higher is closer to $1$.
IV-A1 Program communication
This metric measures the amount of interconnections present in a quantum program, computed as the ratio of the average degree of interaction of the quantum circuit in graph form with that of a maximally connected graph with a number of nodes equal to the number of qubits in the circuit. The term $d(q_{i})$ is the degree of the i-th qubit.
$PC=\frac{\sum_{i}^{N}d(q_{i})}{N(N-1)}$ (6)
IV-A2 Critical depth
The critical depth represents the ratio between the longest chain of two-qubit operators and the total number of two qubit gates in the circuit. It gives a measure on whether the program’s output heavily relies on distributed entanglement or not. $n_{e_{d}}$ is the total number of two qubit gates on the circuit’s critical depth path, while $n_{e}$ is the number of two qubit gates in the circuit.
$CD=n_{e_{d}}/n_{e}$ (7)
IV-A3 Entanglement ratio
This measure is the ratio of the number of entanglement operators, $n_{e}$, with the total number of gates in the quantum circuit, $n_{g}$.
$E=n_{e}/n_{g}$ (8)
IV-A4 Parallelism
It is a measure of the number of concurrent operations made in the same time step, intuitively understood as the degree of compression of the quantum circuit. The number of gates $n_{g}$ is compared with the depth $d$ of the program, then such value is normalised with respect to the number of qubits $n$.
$P=\left(\frac{n_{g}}{d}-1\right)\frac{1}{n-1}$ (9)
IV-B QASMBench
The metrics introduced in the QASMBench [23] suite are more tied to the architectural implementation of physical quantum devices. Follows the definition of the only topologically significant metric.
IV-B1 Entanglement variance
This metric defines the spread of entanglement among the qubits in the quantum circuit. It checks whether there are a few qubits which feature most of connections towards the others, or if all qubits are sharing the same amount of entangling connections. In a quantum program with $N$ qubits, the number of two-qubit gates acting on the i-th qubit is $n_{g_{2}}(q_{i})$, while the average number of two-qubit gates per qubit is $\overline{n_{g_{2}}}$.
$EV=\frac{\log(\sum_{i=0}^{N}(n_{g_{2}}(q_{i})-\overline{n_{g_{2}}})^{2}+1)}{N}$ (10)
TABLE I: The quantum circuits used as benchmarks for the evaluations of this paper.
Circuit name Description Total gates Total multi-qubit gates Ref
QAOA
Quantum Approximate Optimisation Algorithm
$\frac{3}{2}PN(N-1)+2N$
$PN(N-1)$
[13]
Random
Google quantum supremacy circuit
$(1-k)(N(\lfloor N/2\rfloor+N\%2))+kN^{2}$
$kN(\lfloor N/2\rfloor)$
[3]
QPE
Quantum Phase estimation
$\frac{N(N-1)}{2}+2N-1+\lfloor\frac{(N-1)}{2}\rfloor$
$\frac{(N^{2}-N)}{2}+N-2+\lfloor\frac{(N-1)}{2}\rfloor$
[21]
QFT
Quantum Fourier transform
$\frac{1}{2}N(N+1)+\lfloor N/2\rfloor$
$\frac{1}{2}(N^{2}-N)+\lfloor N/2\rfloor$
[10]
VQE
Variational Quantum Eigensolver
$L(5N-1)+N$
$L(N-1)$
[26]
Hamiltonian sim.
One-dimensional Hamiltonian time evolution
$3T(2N-1)$
$T(N-1)$
[31]
Hidden Shift
Find the shift $s$ such that $g(x)=f(x+s)$
$3N+2M+\lfloor N/2\rfloor$
$\lfloor N/2\rfloor$
[34]
Bernstein-Vazirani
Hidden bit string extraction
$2N+M$
$M$
[6]
Figure 5: The scaling of the number of total gates of the circuits considered in this paper, from size 2 to size 32. The Y-axis is in logarithmic scale. The controlling variables are set as $P=1$, $k=0.5$, $M=\lfloor kN\rfloor$, $L=1$, $T=1$. The interactive version of this plot is available online [33].
IV-C Benchmark circuits
In order to provide a broad, extensive and scalable evaluation, we consider a specific set of circuits, selected to encompass some of the applications for quantum computing that do not leverage the presence of quantum noise. As such, their results are significant in terms of exact theoretical simulation. The list of circuits, with information regarding usage, scaling of the number of gates and references, is detailed in Table I. All circuits considered can be freely expanded over any problem size, making them easily adaptable to benchmark future hardware and simulation platforms. For the sake of the hereby presented analysis, we tested all circuit qubit sizes in the range $[2-32]$.
IV-C1 QAOA
The Quantum Approximate Optimisation Algorithm is a variational circuit that use all-to-all connectivity to encode classical problems, such as the max-cut problem. The algorithm version used in this paper is the vanilla one [13] with parameter $P=1$, although other versions exist, such as the one with ZZ-Swap gates [22].
IV-C2 Random
The Random quantum circuit, notably dubbed Quantum Supremacy circuit by the GoogleAI group that introduced it [3], is composed of multiple repeated layers of random gates picked from the set $\mathcal{G}=\{H,X,RZ,RX,RY,CNOT,CZ,SWAP\}$. The number of layers has been set equal to the number of qubits in the system. Given the random nature of the circuit, it is possible to define a lower and an upper bound for the number of gates that can be found in the circuit. This circuit has been purposefully built to avoid any internal structure, so as to be as complex as possible to simulate. Despite achieving such a goal, the applicability for this subroutine to real world problems remains questionable at best.
IV-C3 QPE
The Quantum Phase Estimation subroutine is one of the foundational steps in Shor’s algorithm [21][28], able to solve the order finding problem of a modulo function.
IV-C4 QFT
The Quantum Fourier Transform [10] is one of the most widely known quantum subroutines, which uses phase encoding to efficiently perform the Fourier transform. This quantum circuit is a good candidate for simulation, since it is built by the recursive application of the same operator, possibly giving rise to exploitable internal structures.
IV-C5 VQE
The Variational Quantum Eigensolver is an hybrid quantum-classical algorithm that uses iterative optimisation to find the ground state of a molecule encoded in a quantum register. The circuit is built by repeating L times a given structure. For the sake of simplicity, the version used in this paper assumes $L=1$. Its applications focus mainly on, but are not limited to, the simulation of the bond energies of chemical compounds.
IV-C6 Hamiltonian Simulation
[…]
IV-C7 Hidden Shift
The Hidden Shift quantum circuit is able to find the value $s$ of a function $g(x)=f(x+s)$ by performing a single query to the function, leveraging the superposition of all the possible inputs. The term $k$ represents the percentage of bits equal to 1 in the binary representation of the shift value.
IV-C8 Bernstein-Vazirani
This quantum algorithm can solve the problem of finding out the bit string that satisfies a given function by performing a single query, whilst the classic algorithm would require at most $N$ queries, where $N$ is the total number of possible bit strings. Once again, the term $k$ is the percentage of bits set to 1 in the binary representation of the solution.
V Results
We will first characterise the quantum circuits according to the metrics introduced in Section IV, then we will relate such metrics with the execution time and memory occupancy metrics of each quantum circuit. The library used for running the simulations is the NVIDIA cuQuantum library [29], which has been cross-compiled with the QSim simulator library [30]. This gives us access to three different GPU accelerated simulation backends:
  • qsim-cusv: a state vector simulator that uses the cuStateVec backend.
  • qsim-cuda: a state vector simulator that uses the cupy backend.
  • cutn: a tensor network simulator that uses the cuTensorNet backend for contraction.
All experiments have been run on a system equipped with an AMD EPYC 7643 48-Core Processor, 128 GB of RAM and an NVIDIA A100 80 GB GPU.
Figure 6: Metrics computed for all the benchmark circuits. Given that the topological structure of the Random, Hidden Shift and Bernstein-Vazirani circuits depends on an initialisation seed, the metrics for this circuit have been averaged over 100 different problem instances.
Figure 7: Memory occupancy of a general state vector and the circuit derived tensor networks.
V-A Quantum Circuit Properties
Following the results reported in Figure 6, we will analyse each metric independently. The metrics for the Random, Bernstein-Vazirani and Hidden Shift benchmarks have been averaged over 100 circuit samples, to compensate for the fact that these circuits do not have a constant topology.
The program communication keeps a constant value of $1$ for the QAOA, QPE and QFT circuits, suggesting that those three algorithms feature at least a two-qubit operation with each of the other qubits in the quantum register. This means that the resulting topology of the circuit will be that of a fully connected graph. All other circuits, on the other hand, quickly drop towards values proximal to $0.1$ as soon as the number of qubits in the system increases, meaning that most of the qubits do not interact directly. This can be explained by circuit structures where a small number of qubits interact with all of the others, or by circuit structures where all qubits interact only with their closest neighbours. The main outlier is the random circuit, which stabilises at a value of about $0.5$, meaning that, on average, each qubit interacts with at least half of the total number of qubits in the circuit.
The critical depth starts at value $1$ for most circuits at low qubit sizes and rapidly drops towards the range $[0.08,0.22]$ for the QAOA, QPE, QFT, Random and Hidden Shift benchmarks. This, together with the program communication score, means that the highly entangled structure of the first three circuits is not due to a chain of two qubit operators. The Hidden shift and the Random circuits, having both low program communication and critical depth scores, imply that the derived graph structure is sparsely connected, with a few ”central” qubits sporting most of the two qubit gates towards all other qubits. The remainder of the quantum circuits in the test suite maintain a constant value at $1.0$. This, together with the program communication metric implies that the graph structure of the VQE, Hamiltonian simulation and Bernstein-Vazirani circuits can be reduced to that of a single chain of nodes.
The entanglement ratio attains its maximum value in the QPE and QFT benchmarks, as those circuits are mainly composed of two-qubit gates. The QAOA and Random circuits are composed from $40\%$ to $60\%$ by multiple qubit gates, with the former saturating at $60\%$ as the size of the system increases to 32 qubits, whilst the latter, given the its non deterministic structure, boasts an average of about $50\%$. The other circuits, the Hamiltonian simulation, VQE, Hidden shift and Bernstein-Vazirani, are mainly made of single qubit gates, which can get easily processed during tensor network contraction [16], as such their ER scores are lower, ranging $[0.15-0.35]$.
The parallelism metric grows with respect to the circuit size for all the algorithms considered, with initial values ranging from $0.0$ for the QFT to $0.5$ for the Random. Generally, however, the metric’s value saturates to different levels, with the QAOA, Random, QPE, QFT and Hidden Shift circuits passing the threshold $P>0.8$ for systems sizes of 32 qubits, suggesting that the density of their derived topology is very high. On the other hand, the VQE and Bernstein-Vazirani circuits saturate in the range $[0.6-0.8]$, suggesting a slightly more sparse topology. The Hamiltonian simulation circuit saturates at value $P\approx 0.5$, highlighting its dependence on sequential processing of quantum information and a lower topological density.
Figure 8: At the top, the execution times for the VQE circuit (left) and the Random circuit (right). All data has been collected by doing 3 warmup runs and then averaging the results of 10 additional runs. The tensor network time is the sum of the pathfinding time, done in CPU, and the contraction time, done in GPU. At the bottom, the network structure derived from the same circuits of size 32 qubits, elided of the tensors with degree $d=2$. Blue nodes represent input tensors, while red nodes represent output sensors, node size represents the order of the tensor.
The entanglement variance rapidly approaches zero for almost all circuits considered in the benchmark suite. Notably, the QAOA circuit has a constant variance value of $0.0$, meaning that independently from the circuit size, the number of two qubit operators is evenly split amongst all the qubit in the system. The QPE, QFT, VQE and Hidden Shift algorthms see a rapid decrese in the metric’s value, approaching the range $[0.0-0.05]$ for circuits of 32 qubits, again hinting at the fact that most of the qubits take part in a similar amount of multi-qubit operations. The only two exceptions are the Random and the Bernstein-Vazirani circuits, which instead have a higher value of $ER\in[0.18-0.25]$. In the case of the Random circuit, this is due to the fact that the structire of the circuit does not follow a predefined scheme, whilst in the Bernstein-Vazirani circuit it is directly depended on the number of 1s present in the solution binary bitstring of the oracle function.
V-B Simulation Performance
Figure 7 details the memory requirements scaling for a general state vector simulator and the tensor network representations of all the circuits considered in our benchmark.
The memory occupancy for both the qsim-cuda and the qsim-cusv simulators is the same, as they both have to store in memory the whole state vector of complex probability amplitudes. Each probability amplitude is stored as a complex-64 single precision binary number, scaling exponentially in memory, since the size of the state vector is $N=2^{n}$, with $n$ being the number of qubits in the system. The state vector is updated when applying new quantum gates, however its size remains unaltered, regardless of how many subsequent operations are applied to it.
The tensor network representation instead stores the initial state as a sequence $(1,2)$ tensors, and each quantum gate as either a $(2,2)$ order-2 tensor, in the case of single qubit operators, or as a $(2,2,2,2)$ order-4 tensor, in the case of a controlled gate, adding two dimensions of size 2 for each additional input of the operator. As such, the size of the tensor network representation scales linearly with the number of gates in the quantum circuit and the number of qubits in the system.
In Figure 8, we can see a side by side comparison of the execution times of the VQE and the Random circuits, together with the tensor network structure derived from the circuit, after the removal of all order-2 tensors, for readability’s sake. The time performance data has been collected, for each circuit configuration, as the average time over 10 runs after having performed 3 warmup runs. Thus we collectively performed $3244$ quantum circuit simulations. Code and execution time data relative to these other circuits is available at [2]. We notice how the qsim-cuda simulator has the overall fastest performance for circuits with size of 13 qubits or less, after which the performance of the simulator degrades significantly. This is mainly due to the fact that the shared memory size of the NVIDIA A100, the device we used for simulations, is only configurable up to 164 KB: the state vector of a 13 qubits system, at complex-64 single precision, occupies 64 KB. The performance for circuits of size equal to 14 still holds up, as the state vector needs 128 KB to be stored. However, as soon as the size increases, the L1 cache is not large enough to fit the whole state vector and the simulation gets hindered by memory transfers. From that point onward, the execution time for the qsim-cuda scales exponentially with the system size. Quantum circuits with more than 14 qubits perform better on either of the other two simulators. Notably, the qsim-cusv simulator starts to lose performance for system sizes larger than 22 qubits. This is once again due to the properties of the A100 GPU, which has an L2 cache size of 40 MB, whilst the state vector size of a 22 qubit system is 32 MB. As soon as we increase the system size, the memory requirement doubles and exceeds the cache, forcing additional memory transfers that make the time performance once again scale exponentially in time.
In the case of the Hamiltonian simulation, VQE and QAOA circuits, the tensor network outperforms both state vector simulators. This is a consequence of the fact that these circuits are mainly composed of single qubit gates, corresponding to an entanglement ratio score lower than $0.5$, and have a well distributed amount of entanglement across the system, meaning an entanglement variance score lower than $0.2$. This leads to having small intermediate tensors during the contraction step, that eventually get merged together into larger tensors right towards the end of the process. As soon as one of these two metrics raises up too much, the current pathfinding algorithm [16] cannot find a good enough contraction path.
Figure 9: Heatmap of the speedup of the qsim-cuda and the cutn simulators with respect to the qsim-cuda simulator.
[…]
Moreover, once all the order-2 tensors have been contracted, the resulting topology resembles that of a matrix product state (MPS), which is optimal for contraction. This is an important information, as the contractions of order-2 tensors are easy to perform and do not increase the intermediate tensor order during the contraction process. The circuit with the worst performance for the cutn simulator is the Random circuit, which is the only problem where performance degrades by more than an order of magnitude, mainly due to its lack of an internal structure, something which can be clearly noticed on the right of Figure 8. In all other benchmark circuits, the performance in terms of time is comparable to that of a state vector simulator, whilst keeping the reduced memory footprint of the tensor network representation.
In Figure 9, we can see the relative speedup of the qsim-cusv and the cutn simulators when compared to the qsim-cuda simulator, whose baseline performance is mostly dependent on the system size, rather than on the number of gates to be processed. We can notice negative speedups in the left half of the heatmap, as the qsim-cuda backend outperforms the two other simulators. On the second half of the heatmap, we can appreciate noticeable speedups on both backends. In the QAOA, Random, QPE and QFT circuits the speedups range to up to 60×60$\times$60 × for the cusv backend and up to 53×53$\times$53 × for the cutn backend. We can notice larger speedups on the Hidden Shift, Hamiltonian Simulation, VQE and Bernstein-Vazirani circuits, which exceed the 5000×5000$\times$5000 × speedup value for Hamiltonian Simulation and the VQE. These last four circuits are the ones that scale more slowly in terms of the number of quantum gates, as previously seen in Figure 5. The main outlier is the Random circuit, which has poor performance on the cutn backend, with a negative speedup at high qubit sizes.
Figure 10: The decision diagram to be used for picking the most fitting simulation backend given an arbitrary quantum circuit.
V-C Lessons Learnt
The simulation performance is ruled by a plethora of factors. State vector simulators are inherently limited by the problem size, given the exponential memory requirement, thus trying to to simulate systems larger than $\sim 50$ qubits with state vectors suffers from diminishing returns[38]. Moreover, the distribution of the subvector-matrix multiplications scales exponentially over the problem size, hindering the time performance. The main advantage lies in being able to access the complete set of information encoded in the wave function. Modern state vector simulators can take advantage of GPU caching in order to limit the computation time in small scale problems, but still hit a hard memory bandwidth limit as soon as the cache’s capacity is saturated. Tensor Network contraction has a lot of potential for problems which embody a structure with well balanced connectivity, as that of a structured mesh, as this generates small sized intermediate tensors during the contraction process. As such, current pathfinding algorithms are optimised for finding such structures, and struggle whenever the topology becomes more irregular, such as when most of the entanglement operations in a circuit are concentrated on a few qubits. This is because the size of the intermediate tensors immediately spikes up, slowing the dot product with the remainder of tensors. The main advantage of this approach is the fact that it scales linearly in memory, thus allowing the simulation of larger systems, so long as the circuit’s structure is favourable for contraction. In fact, by scaling the number of qubits in order to saturate the memory capacity of an 80GB NVIDIA A100 GPU, assuming that connectivity information is stored in the shared memory of the machine, one would need more than $7.5k$ logical qubits for a QAOA circuit with $P=1$, or more than $40k$ qubits for the Random circuit, the two circuits in the benchmark with the highest gate scaling per qubit. Despite the fact that there is no guarantee of convergence towards an optimal path, tensor network contraction is the only approach to exact simulation of quantum circuits that can scale favourably to circuits with dimensions larger than $50$ qubits. Previous works show that parallelising the contraction process is trivial, thus the main bottleneck remains the pathfinding algorithm, a well known NP-hard problem [25].
Following the hereby presented results and the statistics collected and discussed in Subsections V-A and V-B, we introduce in Figure 10 a decision diagram to be used for picking the most suitable simulation backend according to the properties of the quantum circuit. As we have observed, circuits which are characterised by an entanglement variance score greater than $0.2$ have a highly unbalanced structure, reducing the efficacy of the pathfinding algorithm for tensor network contraction. Likewise, if more than half of the circuit’s gates are double qubit gates, which amounts to a entanglement ratio score greater than $0.5$, the tensor network approach starts to scale poorly, as stated in Observation 4. This is due to the presence of large tensors early on during the contraction, slowing down the overall process. In both of the previous cases is thus suggested to use a state vector solver. However, if the first two conditions are not met, one must check for the program communication and the critical depth scores. If they are opposite to one another, with one being smaller than $0.2$ and the other being larger than $0.9$, then the best simulation method is the tensor network contraction. This is due to the fact that, if the previous conditions about the entanglement ratio and the entanglement variance scores are met, a program communication score greater than $0.9$ and a critical depth score lower than $0.2$ indicate a circuit structure in which most qubits interact with each other, but there are little to no repeated interactions. On the other hand, a program communication score smaller than $0.15$ together with a critical depth score larger than $0.9$ indicate a circuit structure composed of many chained two-qubit interactions among the same pairs of qubits. In both cases the tensor network becomes a pseudo-regular grid, which can be efficiently contracted whilst keeping intermediate tensor sizes at bay, leading to time performance gains of up to one order of magnitude.
VI Conclusions
The paper characterised the performance of a selected suite of relevant quantum circuits when simulated on state of the art simulator backends and high performance hardware. At first, the circuits have been characterised according to objective metrics that could describe their topological structure, to later correlate them to the performance of the simulators. The results point towards the fact that state vector simulators become heavily communication bound as soon as the size of the statevector exceeds that of the GPU’s cache. The tensor network contraction has proven to perform better in circuits that have well distributed entanglement among the qubits, with a low overall number of two qubit gates with respect to the total number of gates.
It is reasonable to assume that by improving the memory access pattern of a state vector backend, one may improve the time performance of any benchmark to be simulated. On the other hand, the tensor network backend proved to be highly dependent on the efficacy of the pathfinding algorithm. The overall performance of this second backend can already outpace the state vector simulator in some of the benchmark problems, whilst keeping a comparable, albeit slower performance in the remainder of the test runs. It is reasonable to assume that the development of an optimised pathfinding algorithm for these harder to tackle topologies may in fact push the performance of the tensor contraction backend to become the overall fastest simulation approach for quantum circuits. Moreover, implementing a GPU scalable version of the pathfinding algorithm may provide better performance over larger tensor networks without incurring in a computation time trade off. This would indeed be favourable, as tensor network contraction does not suffer from exponential memory requirements as the state vector approach, thus enabling the validation of larger quantum circuits.